# Representation Ensembling for Synergistic Lifelong Learning with Quasilinear Complexity

### v18

@@ Abstract
In lifelong learning, data are used to improve performance not only on the current task, but also on previously encountered, and as yet unencountered tasks. While typical transfer learning algorithms can improve performance on future tasks, their performance on prior tasks degrades upon learning new tasks (called forgetting). Many recent approaches for continual or lifelong learning have attempted to maintain performance on old tasks given new tasks. But striving to avoid forgetting sets the goal unnecessarily low. The goal of lifelong learning should be to use data to improve performance on both future tasks (forward transfer) and past tasks (backward transfer). Our key insight is that we can ensemble representations that were learned independently on disparate tasks to enable both forward and backward transfer, with algorithms that run in quasilinear time. Our algorithms demonstrate both forward and backward transfer in a variety of simulated and benchmark data scenarios, including tabular, vision (CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k), and audition (spoken digit), including adversarial tasks, in contrast to various reference algorithms, which typically failed to transfer either forward or backward, or both.
${}^{1}$ohns Hopkins University, ${}^{2}$aylor College of Medicine, ${}^{3}$niversity of Cambridge, ${}^{4}$icrosoft Research, ${}^{5}$cale AI ${}^{\dagger}$denotes equal contribution, $*{}^{*}$corresponding author: jdey4@jhu.edu
@@ 1 Introduction
In the past 30 years, a number of sequential task learning algorithms have attempted to overcome catastrophic forgetting. These approaches naturally fall into one of two camps. In one, the algorithm has fixed resources, and so must reallocate resources (essentially compressing representations) in order to incorporate new knowledge [10, 11, 12, 13, 14]. For efficient compression of the representation, the model weights can be regularized by using information extracted from old tasks or by directly replaying the old task data. Biologically, this corresponds to adulthood, where brains have a nearly fixed or decreasing number of cells and synapses. In the other, the algorithm adds (or builds) resources as new data arrive (essentially ensembling representations) [15, 16, 17]. Biologically, this corresponds to development, where brains grow by adding cells, synapses, etc. A close resemblance to this resource growing approach can be found in Sodhani et al. [18], where the model adaptively expands when the capacity of the model saturates.
@@
Approaches from both camps demonstrate some degree of continual (or lifelong) learning [19]. In particular, they can sometimes learn new tasks while not catastrophically forgetting old tasks (see Appendix A for a detailed discussion on the relevant algorithms). However, as we will show, many reference lifelong learning algorithms are unable to transfer knowledge forward (to future unseen tasks) and most of them do not transfer backward (to previously seen tasks). With high enough sample sizes, some of them are able to transfer forward or backward, but transfer is more important in low sample size regimes [20, 17]. This inability to effectively transfer in low-sample size regimes has been identified as one of the key obstacles limiting the capabilities of artificial intelligence [21, 22]. We focus primarily on the (arguably simpler) resource growing camp in which each new task is learned with additional representational capacity.
@@
In this paper, We propose two lifelong learning algorithms, one based on ensembling decision forests (Syngeristic Forests, SynF), and another based on ensembling deep networks (Synergistic Networks, SynN). We explore our proposed algorithm as compared to a number of reference algorithms on an extensive suite of numerical experiments that span simulations, vision datasets including CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k, as well as the spoken digit dataset. Figure 1 illustrates that our algorithms outperform all the reference algorithms in terms of forward, backward, and overall transfer in most of the cases. Ablation studies indicate the degree to which the amount of representation or storage capacity and replaying old task data impact performance of our algorithms. All our code and experiments are open source to facilitate reproducibility.
@@
Figure 1: Performance summary on vision and audition benchmark datasets. Columns are datasets, rows are different evaluation criteria (see Section 2 for definitions, and Section 6 for experimental details), each strip of colored dots corresponds to an algorithm (we introduce SynN and SynF here). Colors of dots fade with increasing task number. Datasets on the right have higher number of training samples per task than those in the left. Resource growing algorithms have a ‘*’. EWC, O-EWC, SI, TAG and ER always perform worse than LwF, and hence we do not show them in the summary plot. SynN and SynF tend to exhibit the most transfer across datasets.
@@ 2.1 The lifelong learning objective
Consider a lifelong learning environment with tasks, $\mathcal{T}=\{1,2,\cdots,T\}$. We consider task aware lifelong learning, i.e., the tasks are known during both training and testing time. For simplicity, we consider that any task $t\in\mathcal{T}$ has the same input space, i.e., they have $\mathcal{X}\subset\mathbb{R}^{D}$ valued inputs with $\mathcal{Y}=\{1,\cdots,K_{t}\}$ valued class labels. We assume the tasks arrive sequentially, but the data samples, $\mathbf{s}^{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ within each task $t$ are batched and sampled identically and independently ($iid$) from some fixed distribution. Here each sample within $\mathbf{s}^{t}$ is the realization of random variable pair, $(X,Y)\overset{iid}{\sim}\mathcal{D}_{t}$ and $\mathbf{s}^{t}$ is the realization of $\mathbf{S}^{t}$ distributed as the joint distribution $\mathcal{D}^{n}$. A learner $f\in\mathcal{F}$ trains on $\mathbf{s}^{t}$ and chooses a hypothesis $h\in\mathcal{H}$ by minimizing a particular risk, where $\mathcal{F}$ and $\mathcal{H}$ are the algorithm and hypothesis space, respectively. In supervised learning settings, one can consider the following risk for a particular task $t$:
$R^{t}(f(\mathbf{S}^{t}))=R^{t}(h_{n})=\mathbb{E}_{(X,Y)\sim\mathcal{D}_{t}}[\ell_{t}(h(X),Y)],$ (1)
where $\ell_{t}:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$ is a given loss function associated with the task $t$ and $h=f(\mathbf{S}^{t})$. Note that the data $\mathbf{S}^{t}$ may contain data that is relevant to any number of tasks (potentially all the tasks) in the environment. One may take expectation with respect to $\mathcal{D}^{n}$ for averaging out the randomness in the risk due to $\mathbf{S}^{t}$ and consider the generalization error for the task as:
$\mathcal{E}^{t}_{f}(\mathbf{S}^{t})=\mathbb{E}_{\mathbf{S}^{t}\sim\mathcal{D}^{n}}[R^{t}(f(\mathbf{S}^{t}))].$ (2)
In the above equation, the learner will have access to a total of $T$ datasets after $T$ tasks, $\bigcup_{t=1}^{T}\mathbf{S}^{t}$, instead of $\mathbf{S}^{t}$ only.  [^1] More generally, we may have $J$ datasets, where $J\neq T$ and each dataset may be associated with the target distributions of multiple tasks. For simplicity, we do not consider such scenarios further at this time. The goal is to find a learner $f\in\mathcal{F}$ that chooses a hypothesis $h$ such the generalization error over all the tasks after observing all the data is minimized, that is:
$\begin{array}[]{ll}\textnormal{minimize}&\sum_{t=1}^{T}\mathcal{E}_{f}^{t}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})\\ \textnormal{subject to }&f\in\mathcal{F}\end{array}\enspace.$ (3)
Note that $h$ can operate under any task $t$, i.e., $h=\bigcup_{t=1}^{T}h_{t}$.
@@ 2.2 Lifelong learning evaluation criteria
$\mathsf{Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (4)
@@
$\mathsf{Forward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}.$ (5)
We say an algorithm (positively) forward transfers for task $t$ if and only if $\mathsf{Forward~{}Transfer}^{t}(f)>0$.
Backwards transfer quantifies how much a learner transfers backward to previously observed tasks, in light of new tasks.
@@
$\mathsf{Backward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (6)
We say an algorithm (positively) backward transfers to Task $t$ from all the tasks $T$ if and only if
$\mathsf{Backward~{}Transfer}^{t}(f)>0$.
@@
$\mathsf{Transfer}^{t}(f)$ $=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}+\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}$ (7)
$=\mathsf{Forward~{}Transfer}^{t}(f)+\mathsf{Backward~{}Transfer}^{t}(f).$ (8)
Another paper, Veniat et al. [26], concomitantly introduced transfer and forgetting (backward transfer). Their statistics are the same as ours, except they do not use a $\log$. We opted for a $\log$ to address numerical stability issues in comparing small numbers. Because log is a monotonic function, the order of ranking algorithms is preserved (Appendix Figure 1 shows a version of Figure 1 but using Veniat’s statistics, which is nearly visually identical). By virtue of introducing $\mathsf{Forward~{}Transfer}$ here, we can identify the inherent trade-off between forward and backward transfer, for a fixed amount of total transfer. Apart from the above statistics, we also report accuracy per task.
@@
The accuracy of $f$ on task $t$ is after observing total $T$ datasets is:
$\mathsf{Accuracy}^{t}(f):=1-{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (9)
@@ 3 Representation ensembling algorithms
Figure 2: Schemas of composable hypotheses. A. Single task learner. B. Ensembling decisions (as output by the channels) is a well-established practice, including random forests and gradient boosted trees. C. Learning a joint representation or D. Ensembling representations (learned by the encoders) was previously used in lifelong learning scenarios, but were not trained independently as in E, thereby causing interference or forgetting. Note that the new encoders interact with the previous encoders through the channel layer (indicated by red arrows), thereby, enabling backward transfer. Again the old encoders interact with the future encoders (indicated by black arrows), thereby, enabling forward transfer.
Shannon proposed that a learner can be decomposed into three components: an encoder, a channel, and a decoder [29, 30]: $h(\cdot)=w\circ v\circ u(\cdot)$. Figure 2 shows these three components as the building blocks of different learning schemas. The encoder, $u:\mathcal{X}\mapsto\tilde{\mathcal{X}}$, maps an $\mathcal{X}$-valued input into an internal representation space $\tilde{\mathcal{X}}$ [31, 32]. The channel $v:\tilde{\mathcal{X}}\mapsto\Delta_{\mathcal{Y}}$ maps the transformed data into a posterior distribution (or, more generally, a score). Finally, a decoder $w:\Delta_{\mathcal{Y}}\mapsto\mathcal{Y}$, produces a predicted label.
@@
Our approach also uses multiple encoders and decoders (Figure 2E). Unlike modular approaches, we allow connections across encoders to other channels, including both forwards and backwards. The result is that the channels ensemble representations (learned by the encoders), rather than decisions (learned by the channels). In our algorithms, we push all the data through each encoder, and each channel learns and ensembles across all encoders. When each encoder has learned complementary representations, the channels can leverage that information to improve over single task performance. This approach has applications, particularly in few-shot [39] and multiple task scenarios, including lifelong learning.
@@ 3.1 Our representation ensembling algorithms
We have developed two different representation ensembling algorithms based on bagging which are trained on different tasks. We draw analogy from the synergy of multiple independent representations on a particular task and call our algorithm ‘synergistic’. In both algorithms, as data from a new task arrives, the algorithm first builds a new independent encoder. Then, it builds the channel for this new task by pushing the new task data through all existing encoders. Thus the channel integrates information across all existing encoders using the new task data, thereby enabling forward transfer. At the same time, if it stores old task data (or can generate such data), it can push that data through the new encoders to update the channels from the old tasks, thereby enabling backward transfer. In either case, new test data are passed through all existing encoders and corresponding channels to make a prediction (see Appendix C for detailed description of this approach).
3.1.1 Synergistic Networks
A Synergistic Network (SynN) ensembles deep networks. For each task, the encoder $u_{t}$ in SynN is the “backbone” of a deep network (DN). Thus, each $u_{t}$ maps an element of $\mathcal{X}$ to an element of $\mathbb{R}^{d}$, where $d$ is the number of neurons in the ultimate layer of the DN. The channels are learned via $k$-Nearest Neighbors ($k$-NN) [40] over the $d$ dimensional representations of $\mathcal{X}$. Recall that a $k$-NN, with $k$ chosen such that as the number of samples goes to infinity, $k$ also goes to infinity, while $\frac{k}{n}\to 0$, is a universally consistent classifier [40]. We use $k=16\log_{2}{n}$, which satisfies these conditions. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
3.1.2 Synergistic Forests

### v19

# A Simple Lifelong Learning Approach

@@ Abstract
In lifelong learning, data are used to improve performance not only on the present task, but also on past and future (unencountered) tasks. While typical transfer learning algorithms can improve performance on future tasks, their performance on prior tasks degrades upon learning new tasks (called forgetting). Many recent approaches for continual or lifelong learning have attempted to maintain performance on old tasks given new tasks. But striving to avoid forgetting sets the goal unnecessarily low. The goal of lifelong learning should be to use data to improve performance on both future tasks (forward transfer) and past tasks (backward transfer). In this paper, we show that a simple approach—representation ensembling—demonstrates both forward and backward transfer in a variety of simulated and benchmark data scenarios, including tabular, vision (CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k), and speech (spoken digit), in contrast to various reference algorithms, which typically failed to transfer either forward or backward, or both. Moreover, our proposed approach can flexibly operate with or without a computational budget.
1Johns Hopkins University, 2Baylor College of Medicine, 3University of Cambridge, 4Microsoft Research, denotes equal contribution, corresponding author: jdey4@jhu.edu
@@ 1 Introduction
In the past 30 years, a number of sequential task learning algorithms have attempted to overcome catastrophic forgetting. These approaches naturally fall into one of two camps. In one camp, the algorithm has fixed resources, and so must reallocate resources (essentially compressing representations) in order to incorporate new knowledge [10, 11, 12, 13, 14]. Biologically, this corresponds to adulthood, where brains have a nearly fixed or decreasing number of cells and synapses. In the other camp, the algorithm adds (or builds) resources as new data arrive (essentially ensembling representations) [15, 16, 17, 18]. Biologically, this corresponds to development, where brains grow by adding cells, synapses, etc.
@@
Approaches from both camps demonstrate some degree of continual (or lifelong) learning [19]. In particular, they can sometimes learn new tasks faster due to prior learning on related tasks, while not catastrophically forgetting old tasks (see Appendix .1 for a detailed discussion on the relevant algorithms). However, as we will show, many lifelong learning algorithms are unable to transfer knowledge forward (to future unseen tasks) and most of them do not transfer backward (to previously seen tasks). With high enough sample sizes, some of these algorithms are able to transfer forward or backward, but transfer is more important in low sample size regimes [20, 17]. This inability to effectively transfer in low-sample size regimes has been identified as one of the key obstacles limiting the capabilities of artificial intelligence [21, 22].
@@
In this paper, we propose a general approach for lifelong learning which can be used with many existing encoder models. Specifically, we focus our approach on ensembling deep networks (Simple Lifelong Learning Networks, SiLLy-N). Additionally, we demonstrate how the same approach can be generalized for lifelong learning based on ensembling decision forests (Simple Lifelong Learning Forests, SiLLy-F). We explore our proposed algorithm as compared to a number of reference algorithms on an extensive suite of numerical experiments that span simulation, vision datasets including CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k, as well as the spoken digit dataset. Figure 1 illustrates that our algorithm outperforms all the reference algorithms in terms of forward, backward, and overall transfer on CIFAR 10X10 dataset. Ablation studies indicate the degree to which the amount of representation or storage capacity and replaying old task data impact performance of our algorithms. All our code and experiments are open source to facilitate reproducibility.
@@
Figure 1: Performance summary on CIFAR 10X10 benchmark dataset. Columns are different evaluation criteria (see Section 2 for definitions, and Section 6 for experimental details), each strip of colored dots corresponds to an algorithm (we introduce SiLLy-N here) and each dot represents a task. In all the figure, older tasks have darker colors. Resource growing algorithms have a ‘*’. EWC, O-EWC, SI, TAG and ER always perform worse than LwF, and hence we do not show them in the plot. SiLLy-N (red) outperforms all reference algorithms in terms of forward (second panel), backward (third panel), and overall transfer (first panel). Importantly, such better transfer is achieved at high overall accuracy (last panel). More datasets are evaluated in Figure 8.
@@ 2.1 The lifelong learning objective
Consider a lifelong or continual learning environment with tasks, $\mathcal{T}=\{1,2,\cdots,T\}$. We consider task aware lifelong learning, i.e., the tasks are known during both training and testing time. For simplicity, we consider that any task $t\in\mathcal{T}$ has the same input space, i.e., they have $\mathcal{X}\subset\mathbb{R}^{D}$ valued inputs with $\mathcal{Y}=\{1,\cdots,K_{t}\}$ valued class labels. We assume the tasks arrive sequentially, but the $n_{t}$ data samples within each task $t$, $\mathbf{S}^{t}=\{(X_{i},Y_{i})\}_{i=1}^{n_{t}}$, are batched and sampled identically and independently ($iid$) from some fixed distribution $\mathcal{D}_{t}$, and $\sum_{t=1}^{T}n_{t}=n$. A learner $f\in\mathcal{F}$ trains on $\mathbf{S}^{t}$ and chooses a hypothesis $h_{t}\in\mathcal{H}$ associated with task $t$ by minimizing a particular risk, where $\mathcal{F}$ and $\mathcal{H}$ are the algorithm and the hypothesis space, respectively. In supervised learning settings, one can consider the following risk for a particular task $t$:
(1) $R^{t}(f(\mathbf{S}^{t}))=R^{t}(h_{t})=\mathbb{E}_{(X,Y)\sim\mathcal{D}_{t}}[\ell_{t}(h_{t}(X),Y)],$
where $\ell_{t}:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$ is a given loss function associated with the task $t$ and $h_{t}=f(\mathbf{S}^{t})$. Note that the data $\mathbf{S}^{t}$ may contain data that is relevant to any number of tasks (potentially all the tasks) in the environment. One may take expectation over $\mathbf{S}^{t}$ and consider the generalization error for the task as:
(2) $\mathcal{E}^{t}_{f}(\mathbf{S}^{t})=\mathbb{E}[R^{t}(f(\mathbf{S}^{t}))].$
In the above equation, the learner will have access to a total of $T$ datasets after $T$ tasks, $\bigcup_{t=1}^{T}\mathbf{S}^{t}$, instead of $\mathbf{S}^{t}$ only [^1] More generally, we may have $J$ datasets, where $J\neq T$ and each dataset may be associated with the target distributions of multiple tasks. For simplicity, we do not consider such scenarios further at this time.. The goal is to find a learner $f\in\mathcal{F}$ that chooses a set of total $T$ hypotheses $\{h_{1},\cdots,h_{T}\}$ (one hypothesis for each task) such that the generalization error over all the tasks after observing all the data is minimized, that is:
(3) $\begin{array}[]{ll}\textnormal{minimize}&\sum_{t=1}^{T}\mathcal{E}_{f}^{t}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})\\ \textnormal{subject to }&f\in\mathcal{F}\end{array}\enspace.$
@@ 2.2 Lifelong learning evaluation criteria
(4) $\mathsf{Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
@@
(5) $\mathsf{Forward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}.$
We say an algorithm (positively) forward transfers for task $t$ if and only if:
(6) $\mathsf{Forward~{}Transfer}^{t}(f)>0.$
Backward transfer quantifies how much a learner transfers backward to previously observed tasks, in light of new tasks.
@@
(7) $\mathsf{Backward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
We say an algorithm backward transfers to Task $t$ from all the future tasks up to $T$ if and only if:
(8) $\mathsf{Backward~{}Transfer}^{t}(f)>0.$
@@
(9) $\mathsf{Transfer}^{t}(f)$ $=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}$
(10) $=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}+\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}$
(11) $=$ $\mathsf{Forward~{}Transfer}^{t}(f)+\mathsf{Backward~{}Transfer}^{t}(f).$
Another paper [26], concomitantly introduced transfer and forgetting (backward transfer). Their statistics are the same as ours, except they do not use a $\log$. We opted for a $\log$ to address numerical stability issues in comparing small numbers. Because log is a monotonic function, the order of ranking algorithms is preserved (Appendix Figure 1 shows a version of Figure 1 and Figure 8, but using Veniat’s statistics, which is nearly visually identical). By virtue of introducing $\mathsf{Forward~{}Transfer}$ here, we can identify the inherent trade-off between forward and backward transfer, for a fixed amount of total transfer. Apart from the above statistics, we also report accuracy per task.
@@
The accuracy of algorithm $f$ on task $t$ after observing total $T$ datasets is:
(12) $\mathsf{Accuracy}^{t}(f):=1-{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
@@ 3 Representation ensembling algorithms
Figure 2: Schemas of composable hypotheses. A. Single task learner. B. Ensembling decisions (as output by the channels) is a well-established practice, including random forests and gradient boosted trees. C. Learning a joint representation or D. learning future representations depending on the past encoders was previously used in lifelong learning scenarios, but encoders were not trained independently as in E. Note that the new encoders in E interact with the previous encoders through the channel layer (indicated by red arrows), thereby, enabling backward transfer. Again the old encoders interact with the future encoders (indicated by black arrows), thereby, enabling forward transfer.
Shannon proposed that a learned hypothesis can be decomposed into three components: an encoder, a channel, and a decoder [29, 30]: $h(\cdot)=w\circ v\circ u(\cdot)$. Figure 2 shows these three components as the building blocks of different learning schemas. The encoder, $u:\mathcal{X}\mapsto\tilde{\mathcal{X}}$, maps an $\mathcal{X}$-valued input into an internal representation space $\tilde{\mathcal{X}}$ [31, 32]. The channel $v:\tilde{\mathcal{X}}\mapsto\Delta_{\mathcal{Y}}$ maps the transformed data into a posterior distribution (or, more generally, a score). Finally, a decoder $w:\Delta_{\mathcal{Y}}\mapsto\mathcal{Y}$, produces a predicted label.
@@
Our approach also uses multiple encoders and decoders (Figure 2E). Unlike modular approaches, we allow interaction among encoders through the channels, including both forward and backward interactions. The result is that the channels ensemble representations (learned by the encoders), rather than decisions (learned by the channels as in Figure 2 B). In our algorithms, we push all the data through each encoder, and each channel learns and ensembles across all encoders. When each encoder has learned complementary representations, the channels can leverage that information to improve over single task performance. This approach has applications in few-shot and multiple task scenarios, as well as lifelong learning.
@@ 3.1 Our representation ensembling algorithms
Figure 2E shows a general structure of our algorithm. As data from a new task arrives, the algorithm first builds a new encoder. Then, it builds the channel for this new task by pushing the new task data through all existing encoders. Thus the channel integrates information across all existing encoders using the new task data, thereby enabling forward transfer. At the same time, if it stores old task data (or can generate such data), it can push that data through the new encoders to update the channels from the old tasks, thereby enabling backward transfer. In either case, new test data are passed through all existing encoders and corresponding channels to make a prediction (see appendix for detailed description of this approach).
3.1.1 Simple Lifelong Learning Networks in resource growing mode
A Simple Lifelong Learning Network (SiLLy-N) ensembles deep networks. For each task, the encoder $u_{t}$ in SiLLy-N is the “backbone” of a deep network (DN). Thus, each $u_{t}$ maps an element of $\mathcal{X}$ to an element of $\mathbb{R}^{d}$, where $d$ is the number of neurons in the penultimate layer of the DN. The channels are learned by averaging over $k$-Nearest Neighbors ($k$-NN) [39] trained over the $d$ dimensional representations of $\mathcal{X}$. Note that the channel is trained on the $d$ dimensional outputs from the encoders which is much smaller in size than the original training data and hence, the $k$-NN channels are inexpensive storage-wise (shown later in Figure 3). Other algorithms could also be used to learn the channels, though we do not pursue them here. The decoder $w_{t}$ outputs the $\operatornamewithlimits{argmax}$ to produce a single prediction.
3.1.2 Simple Lifelong Learning Networks in resource constrained mode
The above resource growing approach is ideal when the upcoming tasks become more and more complex and there is no constraint imposed by the computation and storage budget available. However, real-world scenarios often impose computational constraints. In the constant resource mode, we stop building new encoders after we have reached the computation and the storage budget imposed by the user. As new tasks arrive, we only learn new channels associated with new tasks using the old encoders. Note that this approach completely excludes the need to save old task data after we have reached the budget.
Hereafter, we will use suffix ‘-M’ after the algorithm name whenever we use resource constrained operation of SiLLy-N. Here $M$ is the total number of encoder allowed by the budget.
3.1.3 Additional realization of our approach using random forest as encoder